solve step by step
20=500/(10+n)

Answer by macston(5194)

20=500/(10+n) Multiply each side by (10+n)
20(10+n)=500
200+20n=500 Subtract 200 from each side.
20n=300 divide each side by 20.
n=15 ANSWER: n=15